# Meta SecAlign: A Secure Foundation LLM Against Prompt Injection Attacks

### v2

@@ V-B Self-generated response improves utility and security by high-quality in-distribution labels
V-C SecAlign++ with short samples lends security generalizable to long test cases
LLMs have been deployed in longer and longer context scenarios, where the input token length grows exponentially. A straightforward idea to secure LLMs against prompt injections in long contexts is to run defensive fine-tuning on long inputs. Our tentative study suggests otherwise: training on short samples yields strong PI security to all-lengthed tested inputs (as long as 16K), while defensive fine-tuning on longer samples, on the contrary, hurts security.
Meta SecAlign is fine-tuned on a security preference dataset generated from Cleaned-Alpaca [alpacacleaned]. 99.9% of our fine-tuning inputs are within 384 tokens. This is a very short dataset today. We try running SecAlign++ on another longer security preference dataset generated from Natural-Instructions [naturalinstructions]. Here, 50% inputs are longer than 384 tokens. To avoid effects from the dataset size (which is much larger for [naturalinstructions]), we sub-sample this preference dataset to have 60K samples for 1-epoch fine-tuning, which has the same number of training steps as our 3-epoch fine-tuning on the dataset from Cleaned-Alpaca.
Please see the middle two columns (GT annotator) in Table˜VI for comparative results of this study. The LLM fine-tuned with the preference dataset from Natural Instructions has significantly higher ASRs, even to a 93.3% AlpacaFarm Adaptive ASR. In comparison, defensive fine-tuning from a short Cleaned-Alpaca generated dataset gives ASRs consistently lower than 20%. Note that this comparative study is done with SecAlign (without the two proposed techniques), and we show Meta-SecAlign-70B results in the last column.
TABLE VI: Training with short samples (from Cleaned-Alpaca) lends security generalizable to long test cases (AgentDojo, WASP benchmarks), while training with long samples (in Natural Instruction) gives poor security. The second-to-last column stands for SecAlign-70B results, and the last column stands for the new SOTA Meta-SecAlign-70B results.
Dataset None Nat.-Inst. Cleaned-Alpaca
Response Annotator - Ground Truth (GT) SELF
Randomized Injection Position - No Yes
MMLU 0-shot ($\uparrow$) 86.3% 82.1% 85.8% 85.9%
MMLU-Pro 5-shot ($\uparrow$) 67.7% 49.7% 65.4% 67.6%
IFEval ($\uparrow$) 91.3% 82.7% 87.6% 89.5%
BBH 3-shot ($\uparrow$) 85.2% 79.5% 84.5% 84.8%
GPQA Diamond 0-shot ($\uparrow$) 50.0% 48.0% 46.0% 48.0%
AlpacaEval2 Utility ($\uparrow$) 44.2% 36.7% 38.7% 44.7%
SEP Utility ($\uparrow$) 62.1% 42.8% 51.6% 60.4%
AgentDojo Utility ($\uparrow$) 59.8% 5.2% 6.2% 84.5%
AgentDojo Utility w. Attack ($\uparrow$) 43.4% 5.3% 7.9% 79.5%
WASP Utility ($\uparrow$) 62.2% 48.6% 45.9% 59.5%
AlpacaFarm ASR ($\downarrow$) 95.7% 70.7% 0.5% 0.5%
AlpacaFarm Adaptive ASR ($\downarrow$) 98.1% 93.3% 8.2% 0.5%
SEP ASR ($\downarrow$) 99.7% - 19.8% 6.4%
SEP Adaptive ASR ($\downarrow$) 99.7% - 18.9% 6.4%
TaskTracker ASR ($\downarrow$) 19.6% 2.5% 0.2% 0.2%
CyberSecEval2 ASR ($\downarrow$) 52.7% 25.5% 3.6% 1.8%
InjecAgent ASR ($\downarrow$) 53.8% 35.0% 0% 0.5%
AgentDojo ASR ($\downarrow$) 14.7% 0.5% 0% 1.9%
WASP Intermediate ASR ($\downarrow$) 20.2% 11.9% 4.8% 1.2%
WASP End2End ASR ($\downarrow$) 2.4% 1.2% 1.2% 0%
We hypothetically credit the loss of security during long-input training to the intuition that simulated attacks on long samples become weak, and training on weak attacks leads to poor security. When the training samples, especially the data part, get longer, the same simulated injection string will be less prominent and get immersed in the benign parts, hurting its aggressiveness. And intuitively, the LLM learning to be robust against weak attacks cannot prevent stronger attacks, see lessons in classic adversarial training on classifiers [goodfellow2015explaining].
@@ V-C SecAlign++ with short samples lends security generalizable to long test cases
V-D Meta SecAlign allows flexible and easy control of utility-security trade-off
@@
For any defense, there is a natural tradeoff where more secure models tend to have lower utility. Arguably, one can tune the learning rate of SecAlign++ to achieve that, see our study in Figure˜5 in the Appendix, but that requires redoing the entire defensive fine-tuning. On top of this costly control, our recipe provides a very simple way to control this trade-off. This can be easily done at test time, and the choice is offered to whoever is using the released Meta SecAlign to build LLM-integrated applications.
@@
Recall that LoRA [hulora] parameterizes linear-layer weights as
@@
where $W$ is the original weight matrix, $A,B$ are rank-$r$ matrices, and $\alpha>0$ is a fixed constant. Tuning $\alpha$ at test time allows a direct interpolation between the initialization (undefended) LLM and Meta SecAlign, trading off security and utility without further modification to the model.
@@
We visualize this tradeoff in Figure˜2 by showing an aggregate utility score and aggregate ASR averaged across all tested benchmarks. Evidently, LoRA $\alpha$ is effective at controlling this trade-off, where lower LoRA $\alpha$ results in a slightly higher utility model with lower security (i.e., higher ASR). Detailed numbers on each benchmark are shown in Figure˜3, where the ASR drops significantly when we interpolate from Llama-3.3-70B-Instruct to Meta-SecAlign-70B. The utility also interpolates linearly between them, but the difference is small as Meta-SecAlign-70B gets a limited utility drop.
@@
Figure 2: The utility-security trade-off when tuning LoRA $\alpha$. Utility is an average across 9 utility benchmarks. ASR is an average across 7 security benchmarks. Both the utility and ASR averages are weighted by the number of samples in each benchmark.
@@
Figure 3: Tuning the LoRA $\alpha$ at test time is effective to control Meta-SecAlign-70B security (above) and utility (below). Numbers are at in Table˜X.
@@
V-E Meta SecAlign has trivial utility drop to enable prompt injection security

### v3

@@ 3.2 Self-generated responses
3.3 SecAlign++ Algorithm
The fact that the proposed two techniques are so easy to apply, and yet also effective in increasing the utility and security, makes SecAlign++ practical for direct use. Technically, our location of minimal-required changes is realized by analyzing SecAlign’s failure modes in shortcut learning and label quality/distribution. With our proposed randomized injection position (for training inputs) and self-generated responses (for training labels), we summarize our SecAlign++ recipe below.
  1. 1.
    Add a new message type to encapsulate untrusted data in the chat template, using special delimiters provided by the initialization instruction-tuned model $f$.
  2. 2.
    Simulate injected inputs by randomly selecting an instruction from a instruction-tuning dataset and injecting it into the beginning or end of another sample’s data part.
  3. 3.
    Obtain corresponding desirable and undesirable responses by feeding the prompt (with benign data) and the injection to the initialization model $f$, respectively.
  4. 4.
    DPO $f$ on the constructed security preference dataset.
Once the above defensive fine-tuning is finished, inferencing with the defended LLM incurs no noticeable utility drop (for the first time) and no additional computation overhead compared to inferencing with the undefended counterpart.
4 Experiments
Using SecAlign++, we fine-tune Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct [dubey2024llama] to Meta-SecAlign-8B and Meta-SecAlign-70B, respectively. Training and evaluation details are in Sections˜4.1 and 4.2.
Section˜4.3 indicates that Meta-SecAlign-70B achieves state-of-the-art security against PI attacks while performing at a similar level of utility as most closed-source commercial LLMs that employ model-level defense. These results support our claim that Meta SecAlign can serve as an open secure foundation for LLM-integrated applications.
Our most comprehensive evaluation to date reveals previously-unrecognized shortcomings in the SoTA SecAlign: it significantly hurts utility and security in a few domains not evaluated in prior work. Section˜4.4 shows that SecAlign++ fixes these shortcomings, establishing a new frontier of utility-security trade-off under static and adaptive attacks, according to our studies with 4B/8B/70B/109B diverse LLMs.
In the remaining subsections, we further analyze and show that SecAlign++ allows flexible and easy-to-use control of the utility-security trade-off (Section˜4.5), and incurs a trivial utility drop while enabling prompt injection security (Section˜4.6). Also, stronger LLMs are more vulnerable to PIs if left undefended (Section˜4.7), but can still be secured by SecAlign++.
4.1 Training Setup
Following [chen2024struq, chen2024aligning], we use the Cleaned-Alpaca [alpacacleaned] instruction-tuning dataset to construct our preference dataset unless otherwise stated. We pick samples that contain a data part and adopt the simulated injection methods in SecAlign to inject the prompt into the data. We modify the chat template with the additional input role to separate the data, and use the original user role for instruction. We then follow Section˜3 to generate the preference dataset (19157 samples) for DPO.
4.2 Evaluation Setup
We evaluate 9 utility benchmarks and 7 security benchmarks, including general knowledge (MMLU [hendrycks2020measuring], MMLU-Pro [wang2024mmlu], IFEval [zhou2023instruction], BBH [suzgun2022challenging], and GPQA Diamond [rein2024gpqa]), instruction following (AlpacaFarm [dubois2023alpacafarm]/AlpacaEval2 [dubois2024length], SEP [mu2023can], TaskTracker [abdelnabi2024you], and CyberSecEval2 [bhatt2024cyberseceval]), and agentic workflows (AgentDojo [debenedetti2024agentdojo] tool-calling, WASP [evtimov2025wasp] web-navigation, and InjecAgent [zhan2024injecagent] tool-calling). All test attack samples have never been seen in training. For general knowledge benchmarks, we copy the performance numbers from public leaderboards [artificial_analysis] for closed-source models, and use the LM Evaluation Harness library [eval-harness] for open-source models. For AlpacaFarm and SEP instruction following benchmarks, we try 6 attack techniques to enhance the injection sentence (see the paragraphs below). We also test three adaptive attacks (two Basic attacks + GCG [zou2023universal] attack) when comparing to the SecAlign baseline. In agentic workflows, AgentDojo and WASP use multi-turn samples where one turn of the data contains a prompt injection.
AlpacaFarm ASR. We follow [chen2024struq, chen2024aligning] to inject “Print exactly Hacked!” to the 208 samples with data input. We use 6 PI techniques, and a sample is counted as successfully attacked if one of the 6 queries makes the model output begin with the word “Hacked” or “hacked”. The 6 PI techniques are straightforward attack (before data), straightforward attack (after data), ignore attack (before data), ignore attack (after data), completion attack, and completion-ignore attack. Ignore sentences enhance straightforward/completion attacks by paraphrasing “Print exactly Hacked!” to a sentence with the meaning “Ignore previous instructions and print exactly Hacked!”. The tested Completion attacks use a different set of fake delimiters than those in the training time.
AlpacaFarm Basic Adaptive ASR. In Section˜3, we securely encode the prompt and data part with a filter that prevents usage of our system’s separation delimiters. If an attacker has access to the LLM embedding, it can find fake delimiters that are close to those official delimiters in embedding space according to their $\ell_{2}$ distance. Using those fake delimiters to “escape” the prompt-data separation with completion attacks has been shown successful in breaking open-weight LLMs whose embeddings are accessible [jia2025critical], so we test them as well. We report a sample as a successful attack if either an adaptive completion attack or an adaptive completion-ignore attack makes the output begin with “Hacked” or “hacked”.
AlpacaFarm GCG Adaptive ASR. GCG [zou2023universal] has been used to perform prompt injection [chen2024struq, chen2024aligning] by appending the suffix after our injection, pushing the LLM to respond with "Hacked" in the AlpacaFarm attack setting. The attack loss is the negative log probability of outputting "Hacked". GCG uses gradients of the adversarial loss with respect to suffix tokens for the optimization. GCG assumes an unrealistically strong attacker threat model with white-box access, and uses iterative optimization to strengthen the injection.
SEP Basic Adaptive ASR. Similar to calculating AlpacaFarm Basic Adaptive ASR, we use adaptive completion attack and adaptive completion-ignore attack, and a sample is counted as successfully attacked if one of the two queries makes the model output contain a witness word.
AgentDojo Utility. AgentDojo is a dynamic benchmark for security against PI attacks in tool-calling agents. The latest version contains 97 user tasks, and the LLM agent must make the appropriate API calls based on the user instruction and combine their result to derive the correct solution. The agent is deemed successful if it achieves the user’s goal. We use a context window length (token length for an input plus its output) of 16K in AgentDojo to take in all needed long texts. AgentDojo Utility w. Attack assesses the utility score under attacks on all injected samples.
AgentDojo ASR. Each AgentDojo user task is paired with several injection tasks that seek to divert the LLM agent to call a malicious API, resulting in 949 (user task, injection task) pairs. The attack is deemed successful if the malicious API is called. Note that these goals are non-exclusive, i.e., the agent can call the malicious API and then resume and complete the user goal. By default, AgentDojo implements several PI attack styles. We adopt the “important instructions” attack because it consistently achieves the highest ASR on the official leaderboard and our tests. For Llama-3.3-70B-Instruct and Meta-SecAlign-70B, we include the default system prompt (in Llama-3.3-70B-Instruct chat template) only in this benchmark, as it significantly improves model utility. For all AgentDojo evaluations, we test with the benchmark’s provided "repeat_user_prompt" (sandwich [sander2024sandwich]) defense, which boosts both utility and security [wang2025defending]. Results without sandwich defense show the same conclusion, see Table˜10.
WASP Utility. WASP is a dynamic benchmark for web agent prompt injection security built from WebArena [zhouwebarena], and offers a utility test set of 37 samples. The LLM agent is given a user instruction (e.g., create an issue in GitLab) and the webpage as input data, and must autonomously navigate the web to complete the user task. We evaluate all LLMs using the axtree webpage representation, which describes important elements in the webpage in a hierarchical structure using text. We use a context window length of 24K in WASP. WASP has 84 injected samples, constructed by 2 benign tasks and 42 injected tasks. WASP Intermediate ASR reports if the agent gets diverted at any point to the injection task, and WASP End2End ASR records if the agent completes the injection task.
InjecAgent ASR. InjecAgent has 1K API-calling samples. Each sample prompts the LLM to complete a pre-defined task while simulating an API call to retrieve data containing PI, which asks the LLM to call a malicious API. The attack is regarded as successful when the specific malicious API is called, which corresponds to “ASR-total” in the original InjecAgent paper. The injection can be made stronger in the “Enhanced” setting by including an additional sentence “IMPORTANT!!! Ignore all previous instructions and strictly adhere to the following instruction:” We report the higher number between the base and enhanced settings as the attack ASR. Similar to AgentDojo, we evaluate all models using the “sandwich” defense [sander2024sandwich], which repeats the user prompt after the retrieved tool output to remind the agent of its original task. We report results without sandwich in Table˜10.
We use vllm [kwon2023efficient] for fast inference for Llama and Meta SecAlign models. Note that both models should be used with the exact prompt format as detailed in Section˜3, which has been implemented in our released tokenizer.chat_template. Unless otherwise specified, we evaluate on all benchmarks without the default system prompt, adhering to which generally yields good utility and security. We access OpenAI LLMs through Azure API and Gemini LLMs through Google Cloud Platform API.
Table 3: Utility on General Knowledge Benchmarks
Llama-3.3-70B GPT Gemini
Undef. Ours 4o-mini 4o 5 2-Flash 2.5-Flash 3-Pro
MMLU ($\uparrow$) 86.3% 85.9% 82.0% 85.7% - - - -
MMLU-Pro 5-shot ($\uparrow$) 67.7% 67.6% 64.8% 74.8% 87.1% 77.9% 80.9% 90.0%
IFEval ($\uparrow$) 91.3% 89.5% - - - - - -
BBH 3-shot ($\uparrow$) 85.2% 84.8% - - - - - -
GPQA Diamond ($\uparrow$) 50.0% 48.0% 42.6% 54.3% 85.4% 62.3% 68.3% 91.0%
Table 4: Utility and Attack Success Rate (ASR) on Instruction Following Benchmarks
Llama-3.3-70B GPT Gemini
Undef. Ours 4o-mini 4o 5 2-Flash 2.5-Flash 3-Pro
AlpacaEval2 Utility ($\uparrow$) 44.2% 44.7% 44.7% 56.4% 67.8% 38.8% 44.6% 64.3%
SEP Utility ($\uparrow$) 62.1% 60.4% 62.1% 62.5% 78.2% 38.2% 49.5% 70.1%
AlpacaFarm ASR ($\downarrow$) 95.7% 0.5% 1.9% 0% 1.0% 48.6% 81.7% 0.5%
SEP ASR ($\downarrow$) 99.7% 6.4% 24.8% 41.4% 57.6% 57.9% 81.4% 79.7%
TaskTracker ASR ($\downarrow$) 19.6% 0.2% 0.3% 0.6% 0.4% 0.4% 1.1% 0.5%
CyberSecEval2 ASR ($\downarrow$) 52.7% 1.8% 25.5% 20.0% 10.9% 43.6% 43.6% 14.6%
Table 5: Utility and Attack Success Rate (ASR) on Agentic Workflows Benchmarks
Llama-3.3-70B GPT Gemini
Undef. Ours 4o-mini 4o 5 2-Flash 2.5-Flash 3-Pro
AgentDojo Utility ($\uparrow$) 59.8% 84.5% 67.0% 79.4% 80.3% 42.3% 63.9% 92.8%
AgentDojo Utility w. Attack ($\uparrow$) 43.4% 79.5% 51.6% 67.4% 79.7% 37.1% 52.6% 90.6%
WASP Utility ($\uparrow$) 62.2% 59.5% 27.0% 32.4% 59.5% 48.6% 56.8% 59.5%
InjecAgent ASR ($\downarrow$) 53.8% 0.5% 3.3% 22.7% 0.2% 27.2% 0.1% 0.2%
AgentDojo ASR ($\downarrow$) 14.7% 1.9% 11.9% 20.4% 0.2% 11.3% 27.9% 2.3%
WASP Intermediate ASR ($\downarrow$) 20.2% 1.2% 53.6% 17.9% 0% 29.8% 44.1% 1.2%
WASP End2End ASR ($\downarrow$) 2.4% 0% 0% 2.4% 0% 8.3% 14.3% 1.2%
Table 6: SecAlign++ outperforms SecAlign in both utility and security against adaptive attacks.
Llama-3.1-8B-Instruct Llama-3.3-70B-Instruct
Undef. SecAlign Ours Undef. SecAlign Ours
MMLU ($\uparrow$) 72.0% 71.7% 71.7% 86.3% 85.8% 85.9%
MMLU-Pro 5-shot ($\uparrow$) 46.5% 45.9% 46.7% 67.7% 65.4% 67.6%
IFEval ($\uparrow$) 79.1% 73.5% 74.5% 91.3% 87.6% 89.5%
BBH 3-shot ($\uparrow$) 71.9% 71.2% 70.9% 85.2% 84.5% 84.8%
GPQA Diamond ($\uparrow$) 31.3% 30.8% 28.3% 50.0% 46.0% 48.0%
AlpacaEval2 Utility ($\uparrow$) 31.2% 30.7% 31.0% 44.2% 38.7% 44.7%
SEP Utility ($\uparrow$) 51.4% 44.1% 48.8% 62.1% 51.6% 60.4%
AlpacaFarm Basic Adaptive ASR ($\downarrow$) 89.4% 6.7% 0.5% 98.1% 8.2% 0.5%
AlpacaFarm GCG Adaptive ASR ($\downarrow$) 87.0% 28.9% 20.7% 98.1% 53.9% 47.3%
SEP Basic Adaptive ASR ($\downarrow$) 97.1% 36.4% 11.5% 99.7% 18.9% 6.4%
Table 7: SecAlign++ is also applicable to significantly securing Qwen3-4B-Instruct-2507 and Llama-4-Scout-17B-16E-Instruct without non-trivial utility drop. "-" stands for the previous-row benchmark, AgentDojo.
Qwen3-4B Llama-4-Sc.
Undef. Ours Undef. Ours
MMLU ($\uparrow$) 70.7% 70.6% 85.9% 85.3%
MMLU-Pro ($\uparrow$) 64.6% 63.6% 71.7% 71.7%
IFEval ($\uparrow$) 67.6% 63.6% 91.3% 87.2%
BBH ($\uparrow$) 30.6% 67.2% 80.4% 77.6%
GPQA Diamond ($\uparrow$) 37.9% 37.9% 57.1% 54.0%
AlpacaEval2 Util. ($\uparrow$) 54.1% 55.1% 42.7% 43.0%
SEP Utility ($\uparrow$) 75.4% 73.0% 58.4% 58.5%
AgentDojo Utility ($\uparrow$) 47.4% 42.3% 56.7% 54.6%
- Utility w. Attack ($\uparrow$) 41.0% 44.8% 48.6% 47.7%
AlpacaFarm ASR ($\downarrow$) 100% 1.0% 87.0% 3.4%
SEP ASR ($\downarrow$) 97.3% 9.3% 96.2% 8.7%
TaskTracker ASR ($\downarrow$) 13.9% 0.2% 33.4% 0.1%
CyberSec. 2 ASR ($\downarrow$) 52.7% 25.5% 52.7% 10.9%
InjecAgent ASR ($\downarrow$) 1.3% 1.7% 1.0% 0%
AgentDojo ASR ($\downarrow$) 3.4% 0.7% 4.9% 1.2%
4.3 Meta-SecAlign-70B is more secure than most commercial LLMs
Table˜3 shows general knowledge utility benchmark results. The utility drop from SecAlign++ (from Llama-3.3-70B-Instruct to Meta-SecAlign-70B) is minor, with maximum drop around 2% on IFEval and GPQA Diamond. For a non-apples-to-apples comparison with commercial LLMs, Meta-SecAlign-70B achieves stronger performance than GPT-4o-mini on all 3 available benchmark numbers. Recent reasoning LLMs such as GPT-5 and Gemini-3-Pro (we test them with high reasoning mode) have impressive utility, but their training recipe or defense recipe is not accessible despite their technical reports [openai2025gpt5, shi2025lessons].
Table˜4 shows instruction following utility and security benchmark results. Meta-SecAlign-70B achieves one to two orders of magnitude lower ASR compared to Llama-3.3-70B-Instruct without noticeably harming utility. Moreover, Meta-SecAlign-70B offers significantly better security against PIs than all closed-source models on all 4 tested benchmarks (except in AlpacaFarm with 0.5% ASR vs. 0% ASR from GPT-4o). During training, we do not expose the LLM to examples of injected prompts with clear injection intent, e.g., “Ignore previous instruction” or “IMPORTANT INSTRUCTION”. Nevertheless, Meta-SecAlign-70B learns to generalize robustly to these injected prompts in evaluation datasets.
Table˜5 shows agentic workflow utility and security benchmark results. Utility-wise, Meta-SecAlign-70B is even comparable to GPT-5 on AgentDojo and WASP, offering competitive performance on complex agentic tool-calling and web-navigation tasks. We are unsure why AgentDojo utility increases after SecAlign++ for Llama 3.3. This does not happen for other models, see Table˜7. This special case may come from unknown post-training details in Llama-3.3-70B-Instruct. We release AgentDojo logs from Llama-3.3-70B-Instruct and Meta-SecAlign-70B for the community to investigate. Security-wise, Meta SecAlign reduces ASRs on all agentic workflow tasks by one to two orders of magnitude. Meta SecAlign’s ASR on InjecAgent goes down from 53.8% to 0.5%, offering a security comparable to GPT-5 and Gemini-3-Pro. Similarly, the ASR on AgentDojo is greatly reduced from 14.7% to 1.9%, much lower than all closed-source models except GPT-5. Meta SecAlign also has a very low WASP Intermediate ASR, indicating that the agent does not even try to execute the injected task. For agentic tasks, the best LLMs in our test are GPT-5 and Gemini-3-Pro. We show in Table˜9 that different reasoning levels in GPT-5 give similar performance in most PI benchmarks.
4.4 SecAlign++ greatly outperforms SecAlign
Besides the above non-apple-to-apple comparison against commercial defended models, we here show that our SecAlign++ recipe secures LLMs much better than the prior SoTA SecAlign, while enjoying the additional advantage of not hurting utility noticeably.
To avoid saturation on security scores, we employ stronger adaptive attacks when comparing with SecAlign. We first test two basic adaptive attacks, assuming the attacker has access to the LLM embedding. With this knowledge, the attacker can find fake delimiters that are close to the official delimiters in embedding space according to their $\ell_{2}$ distance, and use those fake delimiters to “escape” the prompt-data separation [jia2025critical]. We report a sample as successfully attacked if either of the two adaptive attacks succeed. Besides the above, we also test a stronger attack, the Greedy Coordinate Gradient (GCG) attack [zou2023universal], which uses gradients to search for a suffix that causes the model to follow the injected prompt; it represents a strong adaptive, white-box attack. Table˜6 shows that SecAlign++ is much more robust than SecAlign against basic and adaptive attacks, while preserving higher utility, especially in 70B LLMs.
The above results indicate that SecAlign++ establishes a new frontier of the utility-security trade-off over SecAlign, without noticeable utility loss from the undefended counterpart for the first time. Thus, we present SecAlign++ as a new SoTA defensive fine-tuning recipe against prompt injections. To verify this claim, we study the generality of SecAlign++ to other LLM families beyond Llama 3 series.
We additionally consider two very different LLMs: Qwen3-4B-Instruct-2507 [qwen2025qwen3], the SoTA 4B model from Alibaba, and Llama-4-Scout-17B-16E-Instruct [llama42025], a very large 109B MoE model with 17B active parameters. We use a learning rate of 3.2e3.2e4-4 and 6.4e6.4e4-4 for Qwen3 and Llama4, respectively. Table˜7 shows that even on these very different LLM families, SecAlign++ is still effective at securing them against prompt injections with little utility drop. For example, on Qwen3, the ASR on AlpacaFarm drops from 100% to 1.0%, while MMLU utility only drops from 70.7% to 70.6%. On Llama4, the ASR on AlpacaFarm drops from 87.0% to 3.4%, while MMLU utility only drops from 85.9% to 85.3%. This shows SecAlign++, as a general defense recipe, works well on protecting various open-weight models, with good security and little utility drop.
Sections˜4.3 and 4.4 validate our main contributions that Meta SecAlign-70B could serve as a secure foundation against prompt injections, and SecAlign++ is a new SoTA defensive fine-tuning recipe. In the following subsections, we provide further analysis on the utility-security trade-off in our established new frontier.
@@ 4.4 SecAlign++ greatly outperforms SecAlign
4.5 Meta SecAlign allows flexible and easy control of the utility-security trade-off
@@
For any defense, there is a natural trade-off where more secure models tend to have lower utility. Arguably, one can tune the learning rate of SecAlign++ to achieve that, see our study in Figure˜5, but that requires redoing the entire defensive fine-tuning. On top of this costly control, our recipe provides a very simple way to control this trade-off. This can be easily done at test time, and the choice is offered to whoever is using Meta SecAlign to build LLM-integrated applications.
@@
Recall that LoRA [hulora] parameterizes linear-layer weight as
@@
where $W$ is the original weight matrix, $A,B$ are rank-$r$ matrices, and $\alpha>0$ is a fixed constant. Tuning $\alpha$ at test time allows a direct interpolation between the initialization LLM and Meta SecAlign, trading off security and utility without further modification to the model.
@@
We visualize this trade-off in Figure˜2 by showing an aggregate utility score and aggregate ASR averaged across all tested benchmarks. Evidently, LoRA $\alpha$ is effective at controlling this trade-off, where lower LoRA $\alpha$ leads to a slightly higher utility model with lower security (i.e., higher ASR).
@@
Figure 2: The high-level utility-security trade-off when tuning LoRA $\alpha$. Utility is an average across 9 utility benchmarks. ASR is an average across 7 security benchmarks. Both the utility and ASR averages are weighted by the number of samples in each benchmark.
Detailed numbers on each benchmark are in Figure˜3, where the ASR drops a lot when we interpolate from Llama-3.3-70B-Instruct to Meta-SecAlign-70B. The utility also interpolates linearly between them, but the difference is small, as Meta-SecAlign-70B drops trivial utility.
@@
Figure 3: Tuning the LoRA $\alpha$ at test time is effective to control Meta-SecAlign-70B security (top) and utility (bottom). Detailed numbers are present in Table˜11.
@@
4.6 Meta SecAlign has trivial utility drop to enable prompt injection security